Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Test-Stand, Page 13 of 53  Not logged in ELOG logo
ID Date Author Target/No.# Source Station Module Target Oven W.O# Heat Shield W.O# Status Subject Text Attachments
  807   Tuesday, August 18, 2020, 10:00 Aaron Schmidt       IDP15 Scroll pump installed - RP1 replacedPN: X3815-64000
SN: MY2005SA02
Description: IDP15 Scroll pump, Std.
  
  806   Tuesday, August 11, 2020, 12:04 Aaron SchmidtUCxLP-SIS  44139-1043515-4In progressUCx HS conditioning with PEEK assemblyHS target leg openings were milled. Short
from previous test by the look of it. TBHT
insulator on source side is cracked.
  
  805   Wednesday, August 05, 2020, 10:01 Aaron Schmidt      Problems/SolutionsRoughing pump swap out.Wednesday, August 05, 2020, 10:00: TP6
& TP2 off. IG6 off.
Wednesday, August 05, 2020, 12:20:
  
  804   Tuesday, August 04, 2020, 15:07 Aaron SchmidtUCxLP-SIS  44139-1043515-4In progressUCx HS conditioning with PEEK assembly Tuesday, August 04, 2020, 15:04:
Roughing.

Wednesday, August 05, 2020,
 UCx_-_HS_w-PEEK_Assy_-_43515-4.docx 
  803   Tuesday, July 28, 2020, 13:27 Aaron Schmidt      Problems/SolutionsWater leak on magnet PS Fault: 13334
  
  802   Wednesday, July 22, 2020, 08:56 Luca Egoriti P2N testHP-SISITETM4  In progressp2n heating test - electrical feedthrough configuration The electrical feedthroughs at the
top of the test stand module were reconfigured
to deliver up to 1200A to the p2n target.
 ISACteststand_TMwiring.pptxP2N-TestStand_20200722.pngP2N-TestStand_1200A.png 
  801   Monday, July 20, 2020, 11:40 Aaron Schmidt(Empty-Test)   44160-143822-2In progressConditioning Test P2NSalts added: 
Cs - 15 uL
Li - 10 uL
 P2N_-_1_-_July-20-22-2020.docx 
  800   Monday, July 13, 2020, 14:52 Aaron Schmidt LP-FEBIAD  41790-943515-2 UCx HS conditioning. Broken LP-FEBIAD used to out gas.
Monday, July 13, 2020, 14:52: Roughing.
Monday, July 13, 2020, 15:29: Turbo
  
  799   Monday, July 06, 2020, 05:28 Fernando MaldonadoTest FEBIAD HP-FEBIAD  38295-8 42752-5 DoneFEBIAD development part 3/3Reminder. Mini pump still linked
to SV1. DO NOT OPEN SV2!!!
  
  798   Tuesday, June 30, 2020, 21:59 Fernando MaldonadoTest FEBIADHP-FEBIAD  38295-8 42752-5 DoneFEBIAD development part 2/3 Reminder. Mini pump
still linked to SV1. DO NOT OPEN SV2!!!
  
  797   Thursday, June 25, 2020, 08:57 Fernando MaldonadoTest FEBIADHP-FEBIAD  38295-842752-5DoneFEBIAD development part 1/3 Reminder. Mini
pump still linked to SV1. DO NOT OPEN SV2!!!
  
  796   Monday, June 22, 2020, 07:07 Aaron SchmidtTest FEBIADHP-FEBIAD  38295-842752-5In progressTest FEBIAD (Fernando)Reminder. Mini pump still linked to SV1.
DO NOT OPEN SV2!!!
Monday, June 22, 2020, 08:45: Roughing.
 FEBIAD_Testing_-_June_22_-_23.docxER-580-280-30kV-m28.txt 
  795   Tuesday, June 16, 2020, 11:03 Aaron SchmidtTa#60LP-SISITWTM244139-1343515-5In progressConditioning Ta#60 Tuesday, June 16, 2020, 11:01: Turbo pumps
on.

Wednesday, June 17, 2020,
 Ta#60_Conditioning.docxITW-TM2-Ta#60-LP-SIS_Test_Stand.docxER-580-230-30kV-m8.txtER-580-230-30kV-m40.txt 
  794   Wednesday, June 10, 2020, 08:23 Aaron SchmidtC#1LP-SISITWTM243044-843515-3In progressConditioning C#1Salts: 
Li - 10 uL
Na - 10 uL
 7x 
  793   Tuesday, May 19, 2020, 06:01 Aaron Schmidt     42972-8Problems/SolutionsLeak check / Troubleshoot https://elog.triumf.ca/TIS/Test-Stand/791
 May_19-20_Leak_check_-_FS#3.docxMess_inside_TS_-_1.jpgMess_inside_TS_-_2.jpgBlock_water_leak_visible.jpg 
  792   Tuesday, April 07, 2020, 10:53 Anders Mjos       Vacuum loss / spike / Shutdown


   
        
            
  
  791   Friday, March 27, 2020, 11:30 Aaron Schmidt       Vacuum loss / spike / ShutdownThe ISAC was left with the turbo pumps
and water supplies running. Just the power
supplies off. Monday, March, 23, 2020.
 Mar_27_-_2020_-_Vacuum_Trip.docx 
  790   Tuesday, March 17, 2020, 11:40 Aaron SchmidtSiC#40HP-SISITWTM242218-843515-8In progressConditioning SiC#40Tuesday, March 17, 2020, 11:37: Roughing.
FS# 3 re-connected. All waterlines
re-opened except FS#5 (HS cooling).
 SiC#40_-_Conditioning.docxER-580-230-30kV-m7.txtER-580-230-30kV-m23.txtER-580-230-30kV-m38.txtITW-TM2-SiC#40-HP-SIS_Test_Stand.docx 
  789   Wednesday, March 11, 2020, 13:31 Aaron SchmidtTa#60LP-SIS  44139-1343515-5In progressTa#60 w/ EE testing - 2Thermocouples installed on EE. Double checked
continuity & used a heat gun.
Wednesday, March 11, 2020, 13:30:
 Ta#60_-_TS_Pre-condition_-_EE_Test_-_2.docx 
  788   Wednesday, March 04, 2020, 17:15 Fernando Maldonado      Problems/SolutionsEmittance rig communicationWednesday, March 04, 2020, 17:13: EPICS
page indicated a sucesful scan but data was
empty.
  
ELOG V2.9.2-2455